SV Notes for MJD 56238 (starting evening of November 6)¶
|Observer 1||Observer 2||Run Manager||SISPI Expert on Call||Eyeball Squad||QR|
DECam Configuration and Setup¶
|SISPI Configuration File||sv.ini|
|Image Processing (Calibration)||NCSA|
|Image Processing (First Cut)||NCSA|
Twilight ends 21:37, middle of night to hand over to community SV at 01:26 (but delayed if there is TCS/guider work)
Moonrise at 01:30 58% illuminated, 92 degrees away from our Mini-Survey field on Stripe 82. Dark for our full half-night!
Observing Plan and Goals for the Night¶
- Standard bias and dome flat sequence during afternoon
- Please try to get dome open for temperature equilibration at earliest appropriate time, be sure all cooling/ventilation systems are on.
- Check on DIMM status, running if at all possible.
- Make sure RASICAM is online if weather conditions permit.
- A few times per night ask the operator to contact the other telescopes on Tololo to get their seeing value. Record those in the eLog.
We expect time to be taken tonight to check the new TCS model: ~1-2 hrs
- Photometric standards sequence as soon as dark enough (grizY in 3 fields at different airmasses, preferably one at X < 1.25, one at X = 1.25-1.65 and one at X > 1.65). Follow link for the suggested fields and scripts. Douglas suggests standards can begin 30 min before astronomical twilight.
- In photometric conditions, take BD+17 sequence.
- TCS engineering time
In the case that long exposures are behaving well:
- SN Templates if photometric and seeing <1.2". E1 is pretty well done so skip this and move on to other fields. In the deep fields, focus on g & r first to take advantage of dark night.
- If clear but seeing >1.2", take data for Seeing_Tests
- If thin clouds, any seeing, begin MiniSurvey
- Thicker clouds: DECal sweeps
- Photometric standards before handover time if photometric or thin clouds.
Should we still be unable to take longer exposures reliably:
- VVDS field short exposures, useful for photoz if nothing else
Notes from the 4PM Meeting and Additional Comments¶
Notes from Observers¶
Last night QR processed 201 of 305 exposures (about 65%), with core configuration (10 central CCDs select) and DES default configuration.
Ricardo Ogando was giving support the beginning of the night, I'm pasting some of his notes here:
Exposure 148105 @56238.037 Filter r, ellipticity dropped but after that oscillates
kept like this for 4 exposures, @56238.048 it rised again above e=0.1
Exposure ID: 148119 Filter: r, bad seeing and ellipticity. background seems stable.
FWHM suddenly peaks at ~1.8arcsec for this image and then decreases in the next few exposures till ~1.4arcsec. Ellipticity also reaches ~0.2.
Background seems stable for the sequence of r band exposures. For more details see
Exposure ID: 148128 Filter: r, ellipticity peaks at 0.22
sources are elongated in the N-S direction ?
TASCA b-filter animated GIF for this night (time stamps in UT):